# Aligning Machine and Human Visual Representations across Abstraction Levels

### v2

@@ B.3.1 Principal Components Analysis of Representations
In Fig. 20 we further show projections of the representations using TriMap [Amid and Warmuth, 2019], which is a non-linear dimensionality reduction technique designed to better preserve both the global and the local structure of the data. While the coarse structure is similar to that of the first two PC (see Fig. 18), the fine-grained structure reveals some interesting details. Note for example that in the TriMap projection the ViT-B exhibits a lot of fine-grained structure in the form of tiny clusters.
@@
Figure 20: Projection of the representations of 85k ImageNet images from four different models (columns) onto 2 dimensions using TriMap [Amid and Warmuth, 2019] both before (first row) and after fine-tuning on AligNet (second row). Colors correspond to high-level categories of the images.

### v3

@@ B.3.1 Principal Components Analysis of Representations
In Fig. 21 we further show projections of the representations using TriMap [Amid and Warmuth, 2020], which is a non-linear dimensionality reduction technique designed to better preserve both the global and the local structure of the data. While the coarse structure is similar to that of the first two PCs (see Fig. 19), the fine-grained structure reveals some interesting details. Note for example that in the TriMap projection the ViT-B exhibits a lot of fine-grained structure in the form of tiny clusters.
Figure 21: Projection of the representations of 85k ImageNet images from four different models (columns) onto 2 dimensions using TriMap [Amid and Warmuth, 2020] both before (first row) and after fine-tuning on AligNet (second row). Colors correspond to high-level categories of the images.